OpenStudy (anonymous):

If you have 6,7 what would be the lcm?? I have alot of these I understand them but not really i do but i dont..

OpenStudy (anonymous):

Than makes very little sense... :) Well, 42.

OpenStudy (anonymous):

How would you get 42?? This is all review I just dont understand

OpenStudy (anonymous):

we are looking for the least common multiple. the lowest number that has both 6 and 7 as factors is 42 (the answer to life the universe and everything)

OpenStudy (anonymous):

Ohhh okay thanks! :D

OpenStudy (anonymous):

Dang -_- I broke my pencil

OpenStudy (anonymous):

Have any others? They aren't always that easy.

OpenStudy (anonymous):

Yea 24, 36

OpenStudy (anonymous):

well, been a while since i've done this.

OpenStudy (anonymous):

lets see...

OpenStudy (anonymous):

Yea same for me

OpenStudy (anonymous):

The answer is 72. Now I just need to find how to do that by hand...

OpenStudy (anonymous):

but I think blexting will explain

OpenStudy (anonymous):

24 = 2 * 2 * 2 * 3 36 = 2 * 2 * 3 * 3 Circle numbers they have in common... Put a circle around the 2's and the next 2's and the last 3's .. your lcm is 2 * 2 * 3 (the circled ones) * 2 * 3 the leftover ones. 72

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

Basically if you count by each number 24,48,72,96,120 etc. ( I am counting by 24's) 36,72,108,144,180, etc. (I am counting by 36's) What is the first number they both have.... 72

OpenStudy (anonymous):

Hmm.... I have to let it soak in Lol i get it so all you really have to do is count each number like 24,48... etc and wait until the numbers have the same answer? But whats with the 2's and stuff like that lol

OpenStudy (anonymous):

You can always do the count thing.... you don't have to do the other if you don't want to. it is just quicker.

OpenStudy (anonymous):

Yea its seems quicker...... Way easier to... Thanks!